Carbon tetrachloride with a mass of 1000kg is placed in container with 0.350 min volume. Calculate specific weight.


1
Expert's answer
2020-09-16T21:10:31-0400

The specific weight of carbon tetra-chloride ("\Large CCI\scriptscriptstyle 4" ) is its weight per unit volume.

Mass of "\Large CCI\scriptscriptstyle 4" (m)=1000 kg

volume of "\Large CCI\scriptscriptstyle 4" (v) ="0.350m^{3}" =350 litres.

The weight of "\Large CCI\scriptscriptstyle 4" =m*g

=1000*9.81 ="9810N"

Changing to kilonewtons (kN) ="{9810 \scriptscriptstyle N\over1000\scriptscriptstyle N}*1kN"

w=9.81KN

specific weight = "\frac w V"


="{9.810 \scriptscriptstyle KN\over 0.350\scriptscriptstyle m^3}"

=28.0286 kg/m3